# FastSet: Parallel Claim Settlement

### v1

@@ Abstract
FastSet is an actor-based distributed protocol for decentralized finance and settlement, which is inspired from blockchains. Account holders cooperate by making claims, which can include payments, holding and transferring assets, accessing and updating shared data, medical records, digital identity, and mathematical theorems, among many others. The claims are signed by their owners and are broadcast to a decentralized network of validators, which validate and settle them. Validators replicate the global state of the accounts and need not communicate with each other. In sharp contrast to blockchains, strong consistency is purposely given up as a requirement. Yet, many if not most of the blockchain benefits are preserved. The protocol is proved to be correct, despite its massively parallel nature.
@@ 1 Introduction
A series of papers before 2008 culminating with Bitcoin [16], have incrementally built a belief that a total order on transactions ought to be required in any distributed/decentralized payment system in order to avoid the infamous double spending attack: an account sending two concurrent transactions attempting to spend the same coin with two different recipients. A total ordering enforced on transaction settlement indeed solves the double spending problem, but is it really necessary? Recent works starting with 2019 [8, 9, 3] propose a radically different way to approach the problem, a truly concurrent approach where payments can be generated and settled in different orders by different nodes without breaking the semantics of payments. The key insight of these works is that the order in which an account receives payments is irrelevant, and so is the order in which different accounts send payments—provided each account stays valid: no double spending and sufficient balance. That is, as far as the nodes/validators are in agreement on the set of locally valid transactions that took place in the system, consensus on a total order is unnecessary. We believe that this apparently simple and innocent observation marks a crucial breakthrough moment in decentralized finance. A moment where the tyranny of sequentiality is abolished and the door is open to innovations that will lead to the next generation of decentralized, yet truly concurrent and scalable infrastructure for digital assets.
@@
What makes the problem difficult is that claims can have side effects on the validators’ states. We show that if claims issued by different clients are weakly independent, a notion that generalizes the property of commutativity on payments [8, 9, 3] discussed above to arbitrary state-effectful computations, then the validators’ states will remain (strongly eventually) consistent in spite of the different orders in which they receive and process the claims.
Like in blockchains, FastSet accounts have a state (e.g., a balance) and are required to sign (and pay for) any claims they issue. We present the FastSet protocol abstractly in Section 2, requiring only abstract notions of state, claim, and claim validity and effect. How claims are generated, e.g., randomly by users or programmatically by contracts, is irrelevant for the core protocol and its correctness discussed in Section 3.4. However, to make it practical, implementations of FastSet have to offer specific types of claims and specific ways to generate them. In Section 4 we propose an actor-inspired language for generating claims, which we call the FastSet Language and abbreviate setl. Some accounts are controlled by users, others by setl programs (or scripts, or contracts). The difference is that the user-controlled accounts are free to issue any claims, including ones that create new accounts and interact with them, while the contract-based accounts can only issue claims as prescribed by their setl program/script/contract.
@@ 2.1 Participants
These process claims made by the clients, while at the same time maintain consistency in the global knowledge about the overall system: balances of all tokens and additional data/state of all clients, global state of the entire protocol, like the set of all the claims that were settled, etc.
All participants pocess a key pair consisting of a private signature key and the corresponding public verification key. If msg is a message and $p$ is a participant, then $\langle\textit{msg}\rangle_{p}$ denotes the message signed by $p$, whose authenticity can be publicly checked. Like in blockchains, in FastSet the public key of a client $a$ serves as the public account number, or the address of $a$. Additional layers of privacy can be added if desired, but that is not our concern in this paper.
@@ 2.2 Claims and Weak Independence
Drawing inspiration from concurrency theory (e.g., Mazurkievich traces [15]), where two events are independent iff they can be processed in any order and the final result is the same, we say that two claims $c$ and $c^{\prime}$ are weakly independent, written $c\parallel c^{\prime}$, iff once each of them is independently valid, they can be processed in any order and the final result is the same: $c\parallel c^{\prime}$ iff for any state $s$, if ${c}\!\downarrow_{s}$ and ${c^{\prime}}\!\downarrow_{s}$ then ${c\,c^{\prime}}\!\downarrow_{s}$, ${c^{\prime}\,c}\!\downarrow_{s}$, and $\llbracket{c\,c^{\prime}}\rrbracket{s}=\llbracket{c^{\prime}\,c}\rrbracket{s}$. Therefore, we weaken the classic notion of independence by requiring it to hold only in those states in which both claims are already valid. This assumption is critical for FastSet because it allows to handle payments and asset transfers as particular claims, and thus generalize FastPay [3]. Here are some examples of weakly independent claims:
@@
  • Commutative operations on a shared location: writing the same value, like in setting a one-way flag to ”true”; summing numbers, positive only like in voting or both positive and negative like in reputation systems (e.g., Reddit’s karma); multiplying numbers, like in aggregating signatures; adding elements to a set, like in signing petitions, bidding in an auction, or submitting (an intent to make) a transaction; adding elements to a canonical data-structure, like to a sorted list (which is re-sorted after each addition).
  • Updates of CRDTs, abbreviating Commutative [20, 18] and/or Conflict-free [21] Replicated Data Types, which are data-types used in the context of distributed systems with the property that concurrent updates on them commute. CRDTs were used mainly in concurrent text editing, but the concept is general and includes many interesting examples, such as integer vectors, sets, maps, and even graphs (with some reasonable restrictions).
  • Payments, and more generally, asset transfers from different clients.
@@ 2.3 Assumptions
The validators do not question, nor check the intent behind or the client-specific semantics of the client’s claims. The validators only check the claims’ validity and settle them if valid, operations which are expected to be very efficient; indeed, in practice these operations involve no computations other than updating client balances and storage locations with given values. In other words, from validators’ perspective the client claims are just simple commands that they must comply with after a few sanity checks. Any computations using programs or smart contracts in various programming languages and VMs happen within the clients, using their resources and not validators’. It is not validators’ concern whether client’s computations are adequate or correct according to client’s intended semantics or specs or terms and conditions or whatever promises they may have made to their clients. All the validators need to know is the block of claims that the client issues as a result of those potentially complex computations. As an analogy, the validators play the role of an operating system (OS) on a computer, whose job is to ensure consistency across the various programs being executed; the programs are the clients of the OS and all they do from OS’s perspective is to issue ordered blocks of commands (the claims).
@@
FastSet therefore gives clients ultimate freedom in what claims they can issue. Without any additional verification, it is quite possible that some clients may make mistakes, some of them even maliciously. For example, the client may hold user assets, such as in the case of a bank, or a centralized exchange, or a blockchain. That is, users may send their assets to the client in expectation of some services. Without any additional verification, there is nothing to prevent the client from stealing users’ assets. Even without any evil intent, a client’s account may be operated by a complex program which may have subtle errors (such as unintended non-determinism) and thus sometimes produce an incorrect sequence of claims. Even without any program at all, a user client like Alice may mistype 11 instead of 1 when sending a payment to Bob, in which case additional verification, like a multisig, would have helped.
@@
There is no silver-bullet solution to cover all types of claim verification that clients or applications need or will ever need. To continue to give clients maximum flexibility, yet to allow the honest clients to verifiably prove themselves to their users, FastSet provides the capability for clients to set up verifiers and a quorum among those in order for its claims to be considered valid by the FastSet validators. Verifiers are themselves clients on the network, whose role is to provide specialized verification services. For example, a verifier can be specialized in verifying program executions in EVM (or Python, or Java, etc.) and clients, e.g., blockchains (or AI agents, or exchanges, etc.), can use it. Such a verifier may re-execute the EVM program to check the result, or may verify a proof produced by an external compute, such as a ZK proof or a TEE proof or even just a signature by some trusted authority (e.g., a centralized exchange which offers services to its KYCed clients). It is the client’s full responsibility to decide what verifiers to include in their set and what quorum is sufficient.
@@
To simplify the presentation of FastSet and to uniformly capture more use cases, we assume the existence of a proxy that facilitates the interaction between a client, its verifiers, and the validators. The proxy is just a helper which only affects the liveness, but not the safety of the protocol. All it does is to forward or broadcast messages signed by the various participants, and to blindly aggregate signatures (#$\#\{\textit{sig}_{1},\textit{sig}_{2},...\}$). Should the proxy fail to do its job for whatever reason, anybody can replace it and redo the tasks that were missed, including the client. Proxy can be a client (the sender of the original message containing the claim sequence, or a different client such as a beneficiary or a verifier or a service provider), or a validator, or any combination of these for all or for any of the two services it provides. That is, there could be zero, one, or more proxies for each client or application deployed on FastSet. If zero, then the client initiating the claims can do the job of the proxy, or any of the validators.
@@ 2.4 FastSet
Proxy is a helper and only affects liveness, not safety. It can be a client (the sender of the original message containing the claim sequence, or a different client such as a beneficiary or a verifier or a service provider), a validator, or any combination of these for all or for any of the two services it provides: broadcasting messages to validators and signature aggregation (#$\#\{\textit{sig}_{1},\textit{sig}_{2},...\}$).
1
2
Client $a$ claims $c_{1}\,c_{2}\,...\,c_{k}$ by signing message $m\coloneq\langle c_{1}\,c_{2}\,...\,c_{k},n\rangle_{a}$, where $n$ is its local nonce (incremented after each message). An appropriate set of verifiers $T:=\{t_{1},t_{2},...,t_{l}\}$ approve $m$ by signing it: $\langle m\rangle_{t_{1}}$, $\langle m\rangle_{t_{2}}$, …, $\langle m\rangle_{t_{l}}$. Proxy aggregates the verifiers’ signatures and broadcasts $\langle m\rangle_{\#T}$ to all validators.
3
Upon receiving $\langle m\rangle_{\#T}$ with $m=\langle c_{1}\,c_{2}\,...\,c_{k},n\rangle_{a}$, validator $v$ validates $m$ by checking: signature of $a$ is valid; $v(a).\textit{nonce}=n$; $v(a).\textit{pending}\subseteq\{m\}$; the verifiers’ aggregated signature #$\#T$ is valid and reaches the client-specified quorum; and finally, ${c_{1}\,c_{2}\,...\,c_{k}}\!\downarrow_{v.\textit{state}}$.
4
Validator $v$ signs $m$, sends $\langle m\rangle_{v}$ to proxy, and sets $v(a).\textit{pending}\coloneq\{m\}$.
5
Once validator quorum is reached on message $m$, proxy creates certificate $\langle m\rangle_{\#}$ and broadcasts it to all validators.
6
Upon receival of $\langle m\rangle_{\#}$, each validator $v$: checks the quorum ##\##; checks $m\not\in v.\textit{presettled}\cup v.\textit{settled}$; adds $m$ to $v.\textit{presettled}$. Note that this $m$ in Step 6 can be different from the $m$ in Step 4, even for the same client $a$.
7
Settle any message in $v.\textit{presettle}$ that can be settled: if $m=\langle c_{1}\,c_{2}\,...\,c_{k},n\rangle_{a}$ in $v.\textit{presettled}$ with $v(a).\textit{nonce}=n$ and ${c_{1}\,c_{2}\,...\,c_{k}}\!\downarrow_{v.\textit{state}}$ then $v.\textit{state}\coloneq\llbracket{c_{1}\,c_{2}\,...\,c_{k}}\rrbracket{v.\textit{state}}$, increment $v(a).\textit{nonce}$, reset $v(a).\textit{pending}\coloneq\emptyset$ :- ∅, move $m$ to $v.\textit{settled}$. This step 7 can and should be executed continuously.
@@
Validators’ goal is to consistently replicate a global state, while at the same time to offer services through an API. Each validator $v$ is assumed to maintain the following data at a minimum, and to offer it through its API at request:
@@
Validators will not attempt to validate any received claim block before it accumulates an appropriate number of verifiers which approve it. This information being public, verifiers and clients have multiple ways to synchronize to get this task done. The simplest, but also the least efficient and riskiest for the client, is for the client to just submit its message and assume that its verifiers listen to some validator and will thus get informed when the client message was settled, then verify it, sign it and send it to the proxy which then aggregates all their signatures and sends them to the validators. A more efficient approach is for the client to inform its verifiers as soon as it sends the original message, so they can start their verification process immediately. An even more efficient approach in some applications could be for the client to send its verifiers its intent to execute a program, also known to the verifiers, so they can start their verification while the client is still working on generating its claim block, allowing program execution and its verification to take place in parallel.
@@
Once the validator $v$ has been informed that a client’s claim block $c_{1}\,c_{2}\,...\,c_{k}$ accumulated the quorum of verifiers, it can proceed to validating it. The obvious and cheap checks are the signatures, the nonce, and the verifier quorum. We deliberately left open the decision on where the verifiers and the verifier quorum are stored: some implementations of FastSet may store them in the validator, e.g., $v(a).\textit{verifiers}$ and $v(a).\textit{quorum}$, in which case the check becomes $|T\cap v(a).\textit{verifiers}|\geq v(a).\textit{quorum}$; others may store them as additional items in the message, besides the claim block and nonce; others may include a special claim in the block, say at the end, claiming the verifiers and quorum. In all cases, however, the verifiers and their quorum are part of the contract that the application makes with its users, so it is important that the validators enforce them. The validator also checks that there is no different message sent by the same client with the same nonce that is pending, that is, it checks $v(a).\textit{pending}\subseteq\{m\}$. In other words, if there is any message pending then it must be the exact same message $m$. This ensures that the client does not, intentionally or not, attempt to initiate double-spending attacks or non-deterministic computations. Note that messages from validators can be delayed or lost, so the client or its proxy may attempt to submit $m$ multiple times until they receive validator confirmation, which is why $v(a).\textit{pending}$ is allowed to already contain $m$. Finally, validator $v$ is ready to check the validity of the actual claim block $c_{1}\,c_{2}\,...\,c_{k}$ in its state $v.\textit{state}$, that is, ${c_{1}\,c_{2}\,...\,c_{k}}\!\downarrow_{v.\textit{state}}$. Note that the validator does not update its state at this stage. Indeed, the block has not achieved validator (not verifier) quorum yet, which means that the client may still possibly achieve quorum on a different block with other validators.
@@
When a validator $v$ receives such a valid certificate $\langle m\rangle_{\#}$, it knows with certainty that a quorum of validators have already validated $m$, so it is safe to process and settle $m$. But that needs to be done only once, to avoid applying the same updates twice or more, so $v$ first checks that $m$ is not already settled, or considered to be settled, i.e., presettled. If that is the case, then $v$ should settle $m$. However, $v$ may not be ready to settle $m$, in the sense that the nonce or the claim block of $m$ may not be valid for $v$. In fact, there is no guarantee that $v$ even participated in the quorum of $\langle m\rangle_{\#}$, or that $v$ is even aware of the existence of the client that originally sent $m$, because that particular client and $v$ could have been disconnected until now. Taking into account all these situations, $v$ only adds $m$ to $v.\textit{presettle}$ for now.
@@
The liveness result, Theorem 4 (see also 4 in Theorem 1), tells that if other validators settled more claims than $v$, then $v$ should also be able to make progress and settle more claims. That is, if $v$ continuously scans $v.\textit{presettle}$ it will eventually find messages which are valid and thus can be settled, because FastSet assumes that all messages are eventually delivered. Implementations of FastSet will likely include mechanisms for validators to request updates of missing settled claims from other validators and populate their presettled set with them in order to make faster progress. Note that right before $m$ is settled in this step 7 of the protocol, $v(a).\textit{pending}$ can be either empty or $\{m\}$. The former happens when $v$ did not get a chance to participate in the quorum on $m$, so $v$ never checked ${c_{1}\,c_{2}\,...\,c_{k}}\!\downarrow_{v.\textit{state}}$ (e.g., if $a$ had enough balance to make a payment); however, since $v(a).\textit{nonce}=n$, $v$ is ready to settle $m$ as soon as ${c_{1}\,c_{2}\,...\,c_{k}}\!\downarrow_{v.\textit{state}}$, and it does exactly that. The latter happens when $v$ has already signed $m$ and was waiting for a quorum including other validators as well to be formed. The monotonicity result, Theorem 3 (see also 3 in Theorem 1), tells that it is safe to settle $m$ as is, that is, checking ${c_{1}\,c_{2}\,...\,c_{k}}\!\downarrow_{v.\textit{state}}$ is theoretically unnecessary. However, we choose to keep this definedness check. As seen in Section 4, implementations of FastSet may choose to allow applications to disobey the weak independence requirement. Keeping the definedness check will guarantee that the validator’s state stays well-defined, the damage being contained to locking the non-conforming accounts.
@@
    Determinism The order a validator receives the messages is irrelevant.

### v3

@@ Abstract
FastSet is a distributed protocol for decentralized finance and settlement, which is inspired from both actors and blockchains. Account holders cooperate by making claims, which can include payments, holding and transferring assets, accessing and updating shared data, medical records, digital identity, and mathematical theorems, among others. The claims are signed by their owners and are broadcast to a decentralized network of validators, which validate and settle them. Validators replicate the global state of the accounts and need not communicate with each other. In sharp contrast to blockchains, strong consistency is purposely given up as a requirement. Yet, many if not most of the blockchain benefits are preserved, while capitalizing on actor’s massive parallelism. The protocol is proved to be correct, despite its massively parallel nature.
@@ 1 Introduction
A series of papers before 2008 culminating with Bitcoin [16], have incrementally built a belief that a total order on transactions ought to be required in any distributed/decentralized payment system in order to avoid the infamous double spending attack: an account sending two concurrent transactions attempting to spend the same coin with two different recipients. A total ordering enforced on transaction settlement indeed solves the double spending problem, but is it really necessary? Recent works starting around 2019 [8, 9, 3] propose a radically different way to approach the problem, a truly concurrent approach where payments can be generated and settled in different orders by different nodes without breaking the semantics of payments. The key insight of these works is that the order in which an account receives payments is irrelevant, and so is the order in which different accounts send payments—provided each account stays valid: no double spending and sufficient balance. That is, as far as the nodes/validators are in agreement on the set of locally valid transactions that took place in the system, consensus on a total order is unnecessary. We believe that this apparently simple and innocent observation marks a crucial breakthrough moment in decentralized finance. A moment where the tyranny of sequentiality is abolished and the door is open to innovations that will lead to the next generation of decentralized, yet truly concurrent, safe, and scalable infrastructure for digital assets.
@@
What makes the problem difficult is that claims can have side effects on the validators’ states. However, if claims issued by different clients are weakly independent, a notion that generalizes the property of commutativity on payments [8, 9, 3] discussed above to arbitrary state-effectful computations, then the validators’ states will remain (strongly eventually) consistent in spite of the different orders in which they receive and process the claims.
Like in blockchains, FastSet accounts have a state (e.g., a balance) and are required to sign any claims they issue. We present the protocol in Section 2, requiring only abstract notions of state, claim, and claim validity and effect. How claims are generated, e.g., randomly by users or programmatically by contracts, is irrelevant for the core protocol and its correctness discussed in Section 3.4. However, to make it practical, implementations of FastSet have to offer specific types of claims and specific ways to generate them. In Section 4 we propose an actor-inspired language for generating claims, which we call the FastSet Language and abbreviate setl  [^1] Not to be confused with the SET Language https://en.wikipedia.org/wiki/SETL, invented in the late 1960s, based on the mathematical theory of sets., and pronounce “settle”. Some accounts are controlled by users, others by setl programs (or scripts, or contracts). The difference is that the user-controlled accounts are free to issue any claims, including ones that create new accounts and interact with them, while the contract-based accounts can only issue claims as prescribed by their setl program/script/contract.
@@ 2.1 Participants
These process claims made by the clients, while at the same time maintaining consistency in the global knowledge about the overall system: balances of all tokens and additional data/state of all clients, global state of the entire protocol, like the set of all the claims that were settled, etc.
All participants possess a key pair consisting of a private signature key and the corresponding public verification key. If msg is a message and $p$ is a participant, then $\langle\textit{msg}\rangle_{p}$ denotes the message signed by $p$, whose authenticity can be publicly checked. Like in blockchains, in FastSet the public key of a client $a$ serves as the public account number, or the address of $a$. Additional layers of privacy can be added if desired, but that is out of the scope of this paper.
@@ 2.2 Claims and Weak Independence
Drawing inspiration from concurrency theory (e.g., Mazurkiewicz traces [15]), where two events are independent iff they can be processed in any order with the same result, we say that two claims $c$ and $c^{\prime}$ are weakly independent, written $c\parallel c^{\prime}$, iff once each of them is independently valid, they can be processed in any order and the final result is the same: $c\parallel c^{\prime}$ iff for any state $s$, if ${c}\!\downarrow_{s}$ and ${c^{\prime}}\!\downarrow_{s}$ then ${c\,c^{\prime}}\!\downarrow_{s}$, ${c^{\prime}\,c}\!\downarrow_{s}$, and $\llbracket{c\,c^{\prime}}\rrbracket{s}=\llbracket{c^{\prime}\,c}\rrbracket{s}$. Therefore, we weaken the classic notion of independence by requiring it to hold only in those states in which both claims are already valid. This assumption is critical for FastSet because it allows to handle payments and asset transfers as particular claims, and thus generalize FastPay [3]. Here are some examples of weakly independent claims:
@@
  • Commutative operations on a shared location: writing the same value, like in setting a one-way flag to "true"; adding numbers, positive only like in voting or both positive and negative like in reputation systems (e.g., the karma system of Reddit); multiplying numbers, like in aggregating signatures; adding elements to a set, like in signing petitions, bidding in an auction, or submitting (an intent to make) a transaction; adding elements to a canonical data-structure, like to a sorted list (which is re-sorted after each addition).
  • Updates of CRDTs, abbreviating Commutative [20, 18] or Conflict-free [21] Replicated Data Types, which are data-types used in the context of distributed systems with the property that concurrent updates on them commute. CRDTs were used mainly in concurrent text editing, but the concept is general and includes many interesting examples, such as integer vectors, sets, maps, and even graphs (with some reasonable restrictions).
  • Payments and, more generally, asset transfers from different clients.
@@ 2.3 Assumptions
The validators do not question, nor check the intent behind or the client-specific semantics of the client’s claims. The validators only check the claims’ validity and settle them if valid, operations which are expected to be very efficient; indeed, in practice these operations involve no computations other than updating client balances and storage locations with given values. In other words, from validators’ perspective the client claims are just simple commands that they must comply with after a few sanity checks. Any computations using programs or smart contracts in various programming languages and VMs happen within the clients, using their resources and not validators’. It is not validators’ concern whether client’s computations are adequate or correct according to client’s intended semantics or specs or terms and conditions or whatever promises they may have made to their clients. All the validators need to know is the block of claims that the client issues as a result of those potentially complex computations. As an analogy, the validators play the role of an operating system (OS) in a computer, whose job is to ensure consistency across the various programs being executed; the programs are the clients of the OS and all they do from OS’s perspective is to issue ordered blocks of commands (the claims).
@@
FastSet therefore gives clients ultimate freedom in what claims they can issue. Without any additional verification, it is quite possible that some clients may make mistakes, some of them even maliciously. For example, the client may hold user assets, such as in the case of a bank, or a centralized exchange, or a blockchain. That is, users may send their assets to the client in expectation of some services. Without any additional verification, there is nothing to prevent the client from stealing users’ assets. Even without any evil intent, a client’s account may be operated by a complex program which may have subtle errors (such as unintended non-determinism) and thus sometimes produce an incorrect sequence of claims. Even without any program at all, a user client like Alice may mistype 11 instead of 1 when sending a payment to Bob, in which case additional verification, like a multi-sig, would have helped.
@@
There is no silver-bullet solution to cover all types of claim verification that clients or applications need or will ever need. To continue to give clients maximum flexibility, yet to allow the honest clients to verifiably prove themselves to their users, FastSet provides the capability for clients to set up verifiers and a quorum among those in order for its claims to be considered valid by the FastSet validators. Verifiers are themselves clients on the network, whose role is to provide specialized verification services. For example, a verifier can be specialized in verifying program executions in EVM (or Python, or Java, etc.) and clients, e.g., blockchains (or AI agents, or exchanges, etc.), can use it. Such a verifier may re-execute the EVM program to check the result, or may verify a proof produced by an external compute, such as a ZK proof or a TEE proof or even just a signature by some trusted authority (e.g., a centralized exchange which offers services to its KYC-ed clients). It is the client’s full responsibility to decide what verifiers to include in their set and what quorum is sufficient.
@@
To simplify the presentation of FastSet and to uniformly capture more use cases, we assume the existence of a proxy that facilitates the interaction between a client, its verifiers, and the validators. The proxy is just a helper which only affects the liveness, but not the safety of the protocol. All it does is to forward or broadcast messages signed by the various participants, and to blindly aggregate signatures (#$\#\{\textit{sig}_{1},\textit{sig}_{2},...\}$). Should the proxy fail to do its job for whatever reason, anybody can replace it and redo the tasks that were missed, including the client. Proxy can be a client (the sender of the original message containing the claim sequence, or a different client such as a beneficiary or a verifier or a service provider), or a validator, or any combination of these for all or for any of the two services it provides, namely broadcasting client messages and aggregating validator signatures. That is, there could be zero, one, or more proxies for each client or application deployed on FastSet. If zero, then the client initiating the claims can do the job of the proxy, or any of the validators.
@@ 2.4 FastSet
ClientVerifiersProxyValidators…  …1Claim2Verify4Sign5Confirm3Validate6Presettle7Settle Proxy is a helper and only affects liveness, not safety. It can be a client (the sender of the original message containing the claim sequence, or a different client such as a beneficiary or a verifier or a service provider), a validator, or any combination of these for all or for any of the two services it provides: broadcasting messages to validators and signature aggregation (#$\#\{\textit{sig}_{1},\textit{sig}_{2},...\}$).12 Client $a$ claims $c_{1}\,c_{2}\,...\,c_{k}$ by signing message $m\coloneq\langle c_{1}\,c_{2}\,...\,c_{k},n\rangle_{a}$, where $n$ is its local nonce (incremented after each message). An appropriate set of verifiers $T:=\{t_{1},t_{2},...,t_{l}\}$ approve $m$ by signing it: $\langle m\rangle_{t_{1}}$, $\langle m\rangle_{t_{2}}$, …, $\langle m\rangle_{t_{l}}$. Proxy aggregates the verifiers’ signatures and broadcasts $\langle m\rangle_{\#T}$ to all validators.3 Upon receiving $\langle m\rangle_{\#T}$ with $m=\langle c_{1}\,c_{2}\,...\,c_{k},n\rangle_{a}$, validator $v$ validates $m$ by checking: signature of $a$ is valid; $v(a).\textit{nonce}=n$; $v(a).\textit{pending}\subseteq\{m\}$; the verifiers’ aggregated signature #$\#T$ is valid and reaches the client-specified quorum; and finally, ${c_{1}\,c_{2}\,...\,c_{k}}\!\downarrow_{v.\textit{state}}$.4 Validator $v$ signs $m$, sends $\langle m\rangle_{v}$ to proxy, and sets $v(a).\textit{pending}\coloneq\{m\}$.5 Once a validator quorum is reached on message $m$, proxy creates certificate $\langle m\rangle_{\#}$ and broadcasts it to all validators.6 Upon receiving $\langle m\rangle_{\#}$, each validator $v$: checks the quorum ##\##; checks $m\not\in v.\textit{presettled}\cup v.\textit{settled}$; adds $m$ to $v.\textit{presettled}$. Note that this $m$ in Step 6 can be different from the $m$ in Step 4, even for the same client $a$7 Settle any message in $v.\textit{presettle}$ that can be settled: if $m=\langle c_{1}\,c_{2}\,...\,c_{k},n\rangle_{a}$ in $v.\textit{presettled}$ with $v(a).\textit{nonce}=n$ and ${c_{1}\,c_{2}\,...\,c_{k}}\!\downarrow_{v.\textit{state}}$ then $v.\textit{state}\coloneq\llbracket{c_{1}\,c_{2}\,...\,c_{k}}\rrbracket{v.\textit{state}}$, increment $v(a).\textit{nonce}$, reset $v(a).\textit{pending}\coloneq\emptyset$ :- ∅, move $m$ to $v.\textit{settled}$. This step 7 can and should be executed continuously.
@@
The validator’s goal is to consistently replicate the global state, while at the same time offer services through an API. Each validator $v$ is assumed to maintain the following data at a minimum, and to offer it through its API at request:
@@
Validators will not attempt to validate any received claim block before it accumulates an appropriate number of verifiers which approve it. This information being public, verifiers and clients have multiple ways to synchronize to get this task done. The simplest, but also the least efficient and riskiest for the client, is for the client to just submit its message and assume that its verifiers listen to some validator and will thus get informed when the client message was settled, then verify it, sign it and send it to the proxy which then aggregates all their signatures and sends them to the validators. A more efficient approach is for the client to inform its verifiers as soon as it sends the original message, so they can start their verification process immediately. An even more efficient approach in some applications could be for the client to send its verifiers its intent to execute a program, also known to the verifiers, so that they can start their verification while the client is still working on generating its claim block, allowing program execution and its verification to take place in parallel.
@@
Once the validator $v$ has been informed that a client’s claim block $c_{1}\,c_{2}\,...\,c_{k}$ accumulated the quorum of verifiers, it can proceed to validating it. The obvious and cheap checks are the signatures, the nonce, and the verifier quorum. We have deliberately left open the decision on where the verifiers and the verifier quorum are stored: some implementations of FastSet may store them in the validator, e.g., $v(a).\textit{verifiers}$ and $v(a).\textit{quorum}$, in which case the check becomes $|T\cap v(a).\textit{verifiers}|\geq v(a).\textit{quorum}$; others may store them as additional items in the message, besides the claim block and nonce; others may include a special claim in the block, say at the end, claiming the verifiers and quorum. In all cases, however, the verifiers and their quorum are part of the contract that the application makes with its users, so it is important that the validators enforce them. The validator also checks that there is no different message sent by the same client with the same nonce that is pending, that is, it checks $v(a).\textit{pending}\subseteq\{m\}$. In other words, if there is any message pending then it must be the exact same message $m$. This ensures that the client does not, intentionally or not, attempt to initiate double-spending attacks or non-deterministic computations. Note that messages from validators can be delayed or lost, so the client or its proxy may attempt to submit $m$ multiple times until they receive validator confirmation, which is why $v(a).\textit{pending}$ is allowed to already contain $m$. Finally, validator $v$ is ready to check the validity of the actual claim block $c_{1}\,c_{2}\,...\,c_{k}$ in its state $v.\textit{state}$, that is, ${c_{1}\,c_{2}\,...\,c_{k}}\!\downarrow_{v.\textit{state}}$. Note that the validator does not update its state at this stage. Indeed, the block has not achieved validator (not verifier) quorum yet, which means that the client may still possibly achieve quorum on a different block with other validators.
@@
When a validator $v$ receives such a valid certificate $\langle m\rangle_{\#}$, it knows with certainty that a quorum of validators have already validated $m$, so it is safe to process and settle $m$. But that needs to be done only once, to avoid applying the same updates twice or more, so $v$ first checks that $m$ is not already settled, or considered to be settled, i.e., presettled. If that is the case, then $v$ should settle $m$. However, $v$ may not be ready to settle $m$, in the sense that the nonce or the claim block of $m$ may not be valid for $v$. In fact, there is no guarantee that $v$ even participated in the quorum of $\langle m\rangle_{\#}$, or that $v$ is even aware of the existence of the client that originally sent $m$, because that particular client and $v$ could have been disconnected by now. Taking into account all these situations, $v$ only adds $m$ to $v.\textit{presettle}$ for now.
@@
The liveness result, Theorem 4 (see also 4 in Theorem 1), tells that if other validators settled more claims than $v$, then $v$ should also be able to make progress and settle more claims. That is, if $v$ continuously scans $v.\textit{presettle}$ it will eventually find messages which are valid and thus can be settled, because FastSet assumes that all messages are eventually delivered. Implementations of FastSet will likely include mechanisms for validators to request updates of missing settled claims from other validators and populate their presettled set with them in order to make faster progress. Note that right before $m$ is settled in this step 7 of the protocol, $v(a).\textit{pending}$ can be either empty or $\{m\}$. The former happens when $v$ did not get a chance to participate in the quorum on $m$, so $v$ never checked ${c_{1}\,c_{2}\,...\,c_{k}}\!\downarrow_{v.\textit{state}}$ (e.g., if $a$ had enough balance to make a payment); however, since $v(a).\textit{nonce}=n$, $v$ is ready to settle $m$ as soon as ${c_{1}\,c_{2}\,...\,c_{k}}\!\downarrow_{v.\textit{state}}$, and it does exactly that. The latter happens when $v$ has already signed $m$ and was waiting for a quorum including other validators as well to be formed. The monotonicity result, Theorem 3 (see also 3 in Theorem 1), tells that it is safe to settle $m$ as is, that is, checking ${c_{1}\,c_{2}\,...\,c_{k}}\!\downarrow_{v.\textit{state}}$ is theoretically unnecessary. However, we choose to keep this definedness check. As seen in Section 4, implementations of FastSet may choose to allow applications to disobey the weak independence requirement. Keeping the definedness check will guarantee that the validator’s state stays well-defined, the damage being contained to possibly locking the non-conforming accounts.
@@
    Determinism The order in which a validator receives the messages is irrelevant.